OpenStudy (anonymous):

Which is the formula to calculate area for a circle with diameter eighty-two miles? pi times forty-one squared pi times eighty-two squared two pi times forty-one two pi times eighty-two

OpenStudy (anonymous):

the area of a circles formula is A+pi*r^2 so if your diameter is 82, you want 1/2 of that since radius is 1/2 the diameter so it will be 41 so plug 41 into the equation A=pi*41^2 41 ^2 = 1681 * pi= (using 3.14 as pi, thats how ive always done it) so its approximately 5278 so your anser is A A is the answer, if im correct medal please :D

OpenStudy (anonymous):

thx again

OpenStudy (anonymous):

*anytime